# MaskRIS: Semantic Distortion-aware Data Augmentation for Referring Image Segmentation

### v1

@@ 4.1 Experimental Setups
Datasets. We evaluate our method using three popular benchmarks in RIS: RefCOCO, RefCOCO+, and RefCOCOg. RefCOCO [58] provides 142,209 expressions (3.61 words per expression) for about 50,000 objects across 19,994 images, emphasizing object locations and appearances. RefCOCO+ [58] offers over 141,564 expressions (3.53 words per expression), excluding spatial language to focus on object attributes. RefCOCOg [58] introduces more complexity into the evaluation by including 85,474 longer and more complex expressions (8.43 words per expression) for 54,822 objects in 26,711 images. All datasets are collected on top of the MS COCO dataset [28]. For RefCOCOg, we report results on the UMD partition [58], following the previous studies [49, 31, 23].
Evaluation Metrics. To evaluate model performance, we use mean Intersection-over-Union (mIoU), overall Intersection-over-Union (oIoU), and P@$\mathsf{X}$. While mIoU calculates the average IoU across all test samples, oIoU measures the ratio of the total intersection to the total union areas across all samples. P@$\mathsf{X}$ measures the percentage of test samples with an IoU above a threshold $\mathsf{X}\in\{0.5,0.7,0.9\}$. These metrics ensure a consistent and fair evaluation across different RIS methods.
@@
CRIS [49] RN101 CLIP 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
ETRIS [54] RN101 CLIP 71.06 74.11 66.66 62.23 68.51 52.79 60.28 60.42
RefTR [25] RN101 BERT 74.34 76.77 70.87 66.75 70.58 59.40 66.63 67.39
LAVT [56] Swin-B BERT 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
CGFormer [46] Swin-B BERT 76.93 78.70 73.32 68.56 73.76 61.72 67.57 67.83
@@
LAVT [56] Swin-B BERT 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
CGFormer [46] Swin-B BERT 74.75 77.30 70.64 64.54 71.00 57.14 64.68 65.09
LQMFormer [44] Swin-B BERT 74.16 76.82 71.04 65.91 71.84 57.59 64.73 66.04
NeMo [10] Swin-B BERT 74.48 76.32 71.51 62.86 69.92 55.56 64.40 64.80
ReMamber [55] Mamba-B CLIP 74.54 76.74 70.89 65.00 70.78 57.53 63.90 64.00
CARIS [33] Swin-B BERT 74.67 77.63 71.71 66.17 71.70 57.46 64.66 65.45
@@
PolyFormer [31] Swin-B BERT 74.82 76.64 71.06 67.64 72.89 59.33 67.76 69.05
CARIS [33] Swin-B BERT 76.63 79.40 73.52 68.03 73.70 60.41 67.95 69.75
@@
Implementation Details. To validate MaskRIS, we applied it to various RIS models [54, 49, 56, 33]. Designed as a plug-and-play training strategy, our approach maintains the essential hyperparameters, such as learning rate, epochs, and batch size, as they were in the original settings. Notably, we primarily implemented our method on CARIS [33], a leading SoTA method, unless stated otherwise. Images are resized to 448 $\times$ 448 for both training and testing. For image masking, we set 32 as the patch size.
@@ 4.2 Main Results
Comparison with State of the Arts. We compare MaskRIS with previous methods on three popular benchmarks. As shown in Tab. 1, MaskRIS consistently outperforms previous methods by significant margins. Specifically, compared to the second-best performing method, CARIS [33], our approach improves oIoU scores by 1.82%p, 1.33%p, and 2.25%p on the RefCOCO validation, testA, and testB sets, respectively. On RefCOCO+, our method leads by 1.37%p, 2.76%p, and 1.93%p on the validation, testA, and testB splits, respectively. Even on the challenging RefCOCOg dataset, MaskRIS still outperforms CARIS by 0.89%p and 1.05%p on the validation and test splits. These results validate the effectiveness of our masking strategy for RIS model training.
@@
To further demonstrate the capability of our method, we conduct experiments on a larger, more comprehensive dataset. This dataset is a combination of the training sets from RefCOCO, RefCOCO+, and RefCOCOg, following the approach used in previous studies [31, 33]. Training on this combined dataset yields even better results, with oIoU improvements of 1.17–4.04%p over previous methods. This result indicates that our method is effective across different types of data and highlights its generalizability. Fig. 5 shows qualitative examples on RefCOCO, where our method more comprehensively captures the extent of objects by expanding or reducing their coverage (1st row and 2nd row). Additionally, it demonstrates robustness against occlusion (3rd row) and precisely identifies target objects in alignment with the provided text descriptions (4th row). These results confirm the ability of our approach to enhance model robustness and leverage textual cues effectively.
Figure 6: Robustness on various image corruptions provided by ImageNet-C [12] and linguistically complex situations. The results (oIoU) are evaluated on the RefCOCO validation set.
Robustness Evaluation. A key strength of our approach lies in its robustness across visually and linguistically complex scenarios. To this end, we first evaluate our model’s robustness to various image corruptions. We use the benchmark provided by ImageNet-C [12], which includes 17 types of corruption, each with 5 severity levels. To evaluate the performance, we compute the oIoU at each severity level and average these scores for each corruption type. As shown in Fig. 6(a), MaskRIS consistently outperforms the previous SoTA method, CARIS [33], with oIoU improvements of 1.34–5.07%p. This demonstrates that MaskRIS effectively reduces overfitting to clean data, providing greater robustness against a wide range of visual distortions.
In addition, we evaluate MaskRIS’s robustness to linguistically complex situations, such as occlusion, relative position, and ordering. For occlusion scenarios, we modify the input data by occluding parts of objects. For relative position and ordering, we select samples containing relevant linguistic keywords from the validation set. As shown in Fig. 6(b), MaskRIS achieves superior performance in these complex scenarios on the RefCOCO validation set, further highlighting its robustness to both visual and linguistic complexity.
@@
TRIS [30] RN50 CLIP 31.17 32.43 29.56
TRIS [30]$+$MaskRIS RN50 CLIP 32.60 (+1.43) 34.26 (+1.83) 30.32 (+0.76)
SaG [20] ViT-S/16 BERT 37.21 36.60 39.41
SaG [20]$+$MaskRIS ViT-S/16 BERT 38.85 (+1.64) 37.70 (+1.1) 40.20 (+0.79)
@@
CRIS [49] RN50 CLIP 69.52 72.72 64.70
CRIS [49]$+$MaskRIS RN50 CLIP 70.73 (+1.21) 74.06 (+1.34) 66.82 (+2.12)
ETRIS [54] R101 CLIP 71.06 74.11 66.66
ETRIS [54]$+$MaskRIS R101 CLIP 72.39 (+1.33) 74.99 (+0.88) 68.55 (+1.89)
@@
LAVT [56] Swin-B BERT 72.73 75.82 68.79
LAVT [56]$+$MaskRIS Swin-B BERT 73.79 (+1.06) 76.57 (+0.75) 70.31 (+1.52)
CARIS [33] Swin-B BERT 74.67 77.63 71.71
CARIS [33]$+$MaskRIS Swin-B BERT 76.49 (+1.82) 78.96 (+1.33) 73.96 (+2.25)
@@
Compatibility with Other Methods. Our training strategy is highly compatible, integrates seamlessly with various RIS methods, and achieves significant performance improvements, as shown in Tab. 2. Performance gains range from 0.75%p to 2.25%p, even in challenging scenarios like weakly supervised approaches (e.g., TRIS [30] and SaG [20]), which rely solely on text descriptions without ground truth masks. For TRIS on RefCOCO, our approach increases mIoU by 1.43%p on the validation set, 1.83%p on testA, and 0.76%p on testB. Similarly, our method increases SaG’s performance from 37.21 to 38.85 mIoU on the validation set, from 36.60 to 37.70 on testA, and from 39.41 to 40.20 on testB, achieving improvements of up to 1.64%p. Overall, MaskRIS consistently improves RIS model performance across fully- and weakly-supervised frameworks, demonstrating its adaptability and effectiveness without modifying the original model architecture.
4.3 Ablation Study
Impact of Each Component of MaskRIS. We analyze the impact of each MaskRIS component in model training. Tab. 3 shows segmentation results on the RefCOCO validation set. The results clearly show that masking images and text separately improves model performance. However, the most significant improvement occurs when we apply masking to both images and text together. This demonstrates that the combination of visual and textual masking strategies significantly improves the model’s ability to accurately identify and segment the referred objects. In addition, Distortion-aware Contextual Learning (DCL) provides further performance gains. It effectively mitigates the learning instability caused by input masking, while maintaining the benefits of increased data diversity and robustness.
@@ 4.3 Ablation Study
Fig. 7 illustrates the benefits of image and text masking. CARIS often fails to correctly identify objects when they or their adjacent objects are partially obscured. However, by incorporating image masking, our approach achieves accurate object recognition (1st and 2nd rows). For example, in the scenario shown in the first row with a “baby girl” whose legs are obscured, CARIS inaccurately identifies the legs of an adult as part of the target object. On the other hand, our method successfully separates the obscured elements, demonstrating the superior accuracy of our method in complex environments. Conversely, text masking improves the alignment between the text description and the target object (3rd row). For example, with the “left marine”, CARIS inaccurately identifies the target object by focusing too much on spatial information (i.e., “left”), as shown in Fig. 3(c). Text masking intervenes in such cases, ensuring the accurate recognition of the target object. Furthermore, by integrating both image and text masking, our approach exploits the strengths of each masking strategy.

### v2

@@ 4.1 Experimental Setups
Datasets. We evaluate our method using three popular benchmarks in RIS: RefCOCO, RefCOCO+, and RefCOCOg. RefCOCO (yu2016modeling), all built on the MS COCO dataset (COCO). RefCOCO provides 142,209 expressions for about 50,000 objects across 19,994 images, emphasizing object locations and appearances. RefCOCO+ (yu2016modeling) offers over 141,564 expressions, excluding spatial language to focus on object attributes. RefCOCOg (yu2016modeling) introduces more complexity into the evaluation by including 85,474 longer expressions (8.43 words per expression on average) for 54,822 objects in 26,711 images; for this dataset, we report results on the UMD partition (yu2016modeling), following the previous studies (CRIS; PolyFormer; RISCLIP). To further examine cross-dataset generalization beyond COCO-style images, we additionally evaluate MaskRIS on RefClef, a subset of the ImageCLEF dataset with more diverse natural scenes and object categories (see Appendix A.1 for details).
[…]
@@
CRIS (CRIS) RN101 CLIP 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
ETRIS (ETRIS) RN101 CLIP 71.06 74.11 66.66 62.23 68.51 52.79 60.28 60.42
RefTR (RefTR) RN101 BERT 74.34 76.77 70.87 66.75 70.58 59.40 66.63 67.39
LAVT (LAVT) Swin-B BERT 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
CM-MaskSD (wang2024cm) ViT-L CLIP 74.89 77.54 71.28 67.47 71.80 59.91 66.53 66.63
CGFormer (cgformer) Swin-B BERT 76.93 78.70 73.32 68.56 73.76 61.72 67.57 67.83
@@
LAVT (LAVT) Swin-B BERT 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
CGFormer (cgformer) Swin-B BERT 74.75 77.30 70.64 64.54 71.00 57.14 64.68 65.09
LQMFormer (LQMFormer) Swin-B BERT 74.16 76.82 71.04 65.91 71.84 57.59 64.73 66.04
NeMo (NeMo) Swin-B BERT 74.48 76.32 71.51 62.86 69.92 55.56 64.40 64.80
ReMamber (Remamber) Mamba-B CLIP 74.54 76.74 70.89 65.00 70.78 57.53 63.90 64.00
CARIS (CARIS) Swin-B BERT 74.67 77.63 71.71 66.17 71.70 57.46 64.66 65.45
MagNet (chng2024mask) Swin-B BERT 75.24 78.24 71.05 66.16 71.32 58.14 65.36 66.03
@@
PolyFormer (PolyFormer) Swin-B BERT 74.82 76.64 71.06 67.64 72.89 59.33 67.76 69.05
CARIS (CARIS) Swin-B BERT 76.67 79.85 73.07 68.23 73.98 59.59 67.42 68.73
@@
Implementation Details. To validate MaskRIS, we applied it to various RIS models (ETRIS; CRIS; LAVT; CARIS). Designed as a plug-and-play training strategy, we strictly follow the original training settings and hyperparameters, such as learning rate, epochs, and batch size, without modification. Notably, we primarily implemented our method on CARIS (CARIS), a leading SoTA method, unless stated otherwise. Images are resized to 448 $\times$ 448 for both training and testing. For image masking, we set 32 as the patch size.
@@ 4.2 Main Results
Comparison with State of the Arts. We compare MaskRIS with previous methods on three popular benchmarks. As shown in Table 1, MaskRIS consistently outperforms previous methods by significant margins. Specifically, compared to the second-best performing method, CARIS (CARIS), our approach improves oIoU scores by 1.82%p, 1.33%p, and 2.25%p on the RefCOCO validation, testA, and testB sets, respectively. On RefCOCO+, our method leads by 1.37%p, 2.76%p, and 1.93%p on the validation, testA, and testB splits, respectively. Even on the challenging RefCOCOg dataset, MaskRIS still outperforms CARIS by 0.89%p and 1.05%p on the validation and test splits. These results validate the effectiveness of our masking strategy for RIS model training.
@@
To further demonstrate the capability of our method, we conduct experiments on a larger, more comprehensive dataset. This dataset is a combination of the training sets from RefCOCO, RefCOCO+, and RefCOCOg, following the approach used in previous studies (PolyFormer; CARIS). Training on this combined dataset yields even better results, with oIoU improvements of 1.17–4.04%p over previous methods. This result indicates that our method is effective across different types of data and highlights its generalizability. Figure 6 shows qualitative examples on RefCOCO, where our method more comprehensively captures the extent of objects by expanding or reducing their coverage (1st row and 2nd row). Additionally, it demonstrates robustness against occlusion (3rd row) and precisely identifies target objects in alignment with the provided text descriptions (4th row). These results confirm the ability of our approach to enhance model robustness and leverage textual cues effectively.
Figure 7: Robustness on various image corruptions provided by ImageNet-C (imagenet-c) and linguistically complex situations. The results (oIoU) are evaluated on the RefCOCO validation set.
Robustness Evaluation. A key strength of our approach lies in its robustness across visually and linguistically complex scenarios. To this end, we first evaluate our model’s robustness to various image corruptions. We use the benchmark provided by ImageNet-C (imagenet-c), which includes 17 types of corruption, each with 5 severity levels. To evaluate the performance, we compute the oIoU at each severity level and average these scores for each corruption type. As shown in Figure 7(a), MaskRIS consistently outperforms the previous SoTA method, CARIS (CARIS), with oIoU improvements of 1.34–5.07%p. This demonstrates that MaskRIS effectively reduces overfitting to clean data, providing greater robustness against a wide range of visual distortions.
In addition, we evaluate MaskRIS’s robustness to linguistically complex situations, such as occlusion, relative position, and ordering. For occlusion scenarios, we modify the input data by occluding parts of objects. For relative position and ordering, we select samples containing relevant linguistic keywords from the validation set. As shown in Figure 7(b), MaskRIS achieves superior performance in these complex scenarios on the RefCOCO validation set, further highlighting its robustness to both visual and linguistic complexity.
@@
TRIS (TRIS) RN50 CLIP 31.17 32.43 29.56
TRIS (TRIS)++MaskRIS RN50 CLIP 32.60 (+1.43) 34.26 (+1.83) 30.32 (+0.76)
SaG (SaG) ViT-S/16 BERT 37.21 36.60 39.41
SaG (SaG)++MaskRIS ViT-S/16 BERT 38.85 (+1.64) 37.70 (+1.1) 40.20 (+0.79)
@@
CRIS (CRIS) RN50 CLIP 69.52 72.72 64.70
CRIS (CRIS)++MaskRIS RN50 CLIP 70.73 (+1.21) 74.06 (+1.34) 66.82 (+2.12)
ETRIS (ETRIS) R101 CLIP 71.06 74.11 66.66
ETRIS (ETRIS)++MaskRIS R101 CLIP 72.39 (+1.33) 74.99 (+0.88) 68.55 (+1.89)
@@
LAVT (LAVT) Swin-B BERT 72.73 75.82 68.79
LAVT (LAVT)++MaskRIS Swin-B BERT 73.79 (+1.06) 76.57 (+0.75) 70.31 (+1.52)
ReMamber (Remamber) Mamba-B CLIP 72.37 74.85 68.29
ReMamber (Remamber)++MaskRIS Mamba-B CLIP 74.02 (+1.65) 76.32 (+1.47) 69.06 (+0.77)
DETRIS (DETRIS) DINOv2-B CLIP 74.47 77.40 71.46
DETRIS (DETRIS)++MaskRIS DINOv2-B CLIP 76.08 (+1.61) 78.46 (+1.06) 72.84 (+1.38)
CARIS (CARIS) Swin-B BERT 74.67 77.63 71.71
CARIS (CARIS)++MaskRIS Swin-B BERT 76.49 (+1.82) 78.96 (+1.33) 73.96 (+2.25)
@@
Compatibility with Other Methods. Our training strategy is highly compatible, integrates seamlessly with various RIS methods, and achieves significant performance improvements across both weakly supervised and fully supervised settings, as shown in Table 2. In the weakly supervised regime, where models rely solely on textual descriptions without ground-truth masks, MaskRIS still yields notable gains. For instance, TRIS (TRIS) improves by +1.43%p on the validation set, +1.83%p on testA, and +0.76%p on testB. Likewise, SaG (SaG) is enhanced by up to +1.64%p, reaching 38.85 mIoU on the validation set. In the fully supervised setting, MaskRIS demonstrates equally strong compatibility. CRIS (CRIS) shows improvements of +1.21%p (val), +1.34%p (testA), and +2.12%p (testB), while ETRIS (ETRIS) gains up to +1.89%p. Beyond CNN-based baselines, transformer-based approaches such as LAVT (LAVT) and CARIS (CARIS) also benefit significantly, with CARIS achieving the highest improvements of +2.25%p on testB. More recent architectures, including ReMamber (Remamber) and DETRIS (DETRIS), further validate the generality of our method, consistently obtaining +0.8–2.0%p gains without architectural modifications. Overall, MaskRIS consistently improves RIS model performance across diverse architectures and training paradigms, demonstrating its adaptability and effectiveness without modifying the original designs.
Method Image Encoder Text Encoder RefCOCO
val testA testB
DetAcc
EEVG (EEVG) ViT-B BERT 87.87 89.97 85.34
EEVG (EEVG)++MaskRIS ViT-B BERT 88.72 (+0.85) 90.81 (+0.84) 85.36 (+0.02)
SimVG (SimVG) ViT-B BEiT-3 86.94 88.97 82.87
SimVG (SimVG)++MaskRIS ViT-B BEiT-3 87.79 (+0.85) 90.02 (+1.05) 84.88 (+2.01)
C3VG (C3VG) ViT-B BEiT-3 92.42 94.12 89.34
C3VG (C3VG)++MaskRIS ViT-B BEiT-3 93.29 (+0.87) 94.86 (+0.74) 90.29 (+0.95)
Table 3: Compatibility of MaskRIS with REC methods. MaskRIS consistently enhances existing methods.
Compatibility with REC Methods. The conflict between conventional image augmentations and language-grounded tasks is not unique to RIS. Referring Expression Comprehension (REC) also relies on precise spatial and attribute alignment between text and image. For example, flipping an image containing the phrase “the man on the left” alters the semantic meaning for both RIS and REC. This confirms that the augmentation dilemma is indeed universal to referring tasks. To investigate this, we integrated MaskRIS into various REC methods (EEVG; SimVG; C3VG). As shown in Table 3, MaskRIS consistently improves performance when combined with existing REC models by enhancing robustness to occlusion and incomplete linguistic cues while increasing data diversity. The findings confirm that MaskRIS is not task-specific. Its masking approach addresses the augmentation dilemma across both RIS and REC, making it a general training strategy for referring tasks where traditional augmentations are unsafe.
4.3 Ablation Study
Impact of Each Component of MaskRIS. We analyze the impact of each MaskRIS component in model training. Table 4 shows segmentation results on the RefCOCO validation set. The results clearly show that masking images and text separately improves model performance. However, the most significant improvement occurs when we apply masking to both images and text together. This demonstrates that the combination of visual and textual masking strategies significantly improves the model’s ability to accurately identify and segment the referred objects. In addition, Distortion-aware Contextual Learning (DCL) provides further performance gains. It effectively mitigates the learning instability caused by input masking, while maintaining the benefits of increased data diversity and robustness.
@@ 4.3 Ablation Study
Figure 8 illustrates the benefits of image and text masking. CARIS often fails to correctly identify objects when they or their adjacent objects are partially obscured. However, by incorporating image masking, our approach achieves accurate object recognition (1st and 2nd rows). For example, in the scenario shown in the first row with a “baby girl” whose legs are obscured, CARIS inaccurately identifies the legs of an adult as part of the target object. On the other hand, our method successfully separates the obscured elements, demonstrating the superior accuracy of our method in complex environments. Conversely, text masking improves the alignment between the text description and the target object (3rd row). For example, with the “left marine”, CARIS inaccurately identifies the target object by focusing too much on spatial information (i.e., “left”), as shown in Figure 4(c). Text masking intervenes in such cases, ensuring the accurate recognition of the target object. Furthermore, by integrating both image and text masking, our approach exploits the strengths of each masking strategy.
4.4 Discussion: Comparison with CM-MaskSD
Masking Image Enc. Text Enc. oIoU Object-centric Swin-B BERT 75.14 Context 75.93 Holistic 76.02
Table 5: Comparison of image masking approaches on the RefCOCO validation set.
MaskRIS differs from CM-MaskSD (wang2024cm) in three key aspects: purpose, masking strategy, and adaptability. First, MaskRIS explicitly handles the semantic conflicts caused by conventional augmentations in RIS, ensuring that transformations do not alter the contextual meaning of referring expressions. In contrast, CM-MaskSD focuses on improving patch-text alignment using CLIP’s global vector but does not directly tackle augmentation-induced inconsistencies. This makes semantic conflict resolution a unique contribution of MaskRIS. Second, while CM-MaskSD employs object-centric masking, selectively targeting key objects or keywords, MaskRIS utilizes a holistic masking approach that applies masking across all image regions. This strategy enhances model generalization and robustness. As shown in Table 5, holistic masking consistently outperforms object-centric masking. Third, CM-MaskSD is architecture-dependent due to its reliance on the [CLS] token, limiting its applicability to CLIP-based models. In contrast, MaskRIS is architecture-agnostic and can be seamlessly integrated into various RIS and REC frameworks.
5 Conclusion
Acknowledgments
Some parts of experiments are based on the NAVER Smart Machine Learning NSML (kim2018nsml) platform. This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02217259, 25%), the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (RS-2025-00520207, 25%, RS-2023-00219019, 25%), and Samsung Electronics Co., Ltd (IO230508-06190-01, 25%).
Supplementary Material
Contents
Appendix A Additional Experiments
Method
Image
Encoder
Text
Encoder
oIoU
CARIS (CARIS) Swin-B BERT 75.21
CARIS + MaskRIS 76.06
Table A: Comparison on RefClef (ImageCLEF) dataset. MaskRIS consistently improves over the baseline, demonstrating generalization beyond MS COCO benchmarks.
A.1 Evaluation on RefClef
Following the common practice of prior works in RIS, our main experiments are conducted on MS COCO–based benchmarks (RefCOCO, RefCOCO+, RefCOCOg). In addition, we validated MaskRIS on RefClef (kazemzadeh2014referitgame), which is a subset of the ImageCLEF dataset. Unlike MS COCO, RefClef includes more diverse natural scenes and a broader set of object categories beyond MS-COCO’s predefined classes.
As shown in Table A, MaskRIS outperforms the CARIS baseline (76.06 vs. 75.21), confirming that its effectiveness is not limited to COCO-style data. These results demonstrate that our holistic context masking and distortion-aware training framework generalizes well across datasets with varying scene compositions and category coverage, thereby reinforcing the broader applicability of MaskRIS.
Figure A: Various types of image mask strategies. We use patch-wise sampling (MAE) as our default. For block-wise sampling, we follow BEiT(BEIT) to remove large random blocks. For Cutout (CUTOUT), we follow the implementation of timm (timm). For all strategies except Cutout, we maintain a consistent masking ratio of 75%.
Patch-wise Grid-wise Block-wise Cutout
oIoU 76.49 76.14 75.90 76.07
Table B: Impact of imaeg masking strategy. The performance (oIoU) is evaluated across various masking strategies on the RefCOCO validation set.
MaskRIS WWM SpanBERT POS-based
oIoU 76.49 76.03 76.46 76.52
Table C: Impact of text masking strategy. The performance (oIoU) is evaluated across various masking strategies on the RefCOCO validation set.
A.2 Analysis of Image Masking Strategy
To rigorously validate the impact of image masking on model training, we explored various mask sampling strategies and their effects. Figure A visually illustrates each method. Our baseline, patch-wise sampling (MaskSub; MIC; MAE) divides images into non-overlapping patches and randomly masks certain patches. This approach is widely used due to its simplicity and effectiveness. We also tested grid-wise sampling (MAE), where rows and columns are masked systematically rather than randomly. Additionally, we investigated block-wise masking, as used in BEiT (BEIT), where contiguous regions of the image are masked, potentially eliminating more important structural details by covering larger areas. Finally, we evaluated Cutout (CUTOUT), which masks a randomly selected rectangular section of the image.
Table B summarizes the results (oIoU) of each mask sampling strategy on the RefCOCO validation set. Remarkably, all the strategies we tested consistently outperform the previous SoTA method, CARIS (CARIS) (74.67 oIoU), with significant margins. Among these strategies, our patch-wise random sampling approach achieved superior performance, demonstrating its ability to generate more varied training data than grid-wise sampling and minimize training data distortion more effectively than block-wise sampling. These results confirm the importance of selecting an appropriate masking strategy to improve performance.
A.3 Analysis of Text Masking Strategy
To further validate the role of linguistic perturbation, we conducted an ablation study on various text masking strategies. Specifically, we compared four schemes. (1) Token-level random masking, which is adopted in MaskRIS. (2) Whole-Word Masking (WWM) (WWM), which masks all sub-tokens of a sampled word (e.g., masking “playing” removes both “play” and “##ing”), often hiding entire entities or key terms. (3) SpanBERT-style masking (SpanBERT), which replaces contiguous spans of tokens (e.g., masking a two-word phrase like “on the”). This encourages the model to capture longer-range contextual dependencies. (4) Part-of-Speech (POS)-aware random masking, where the target POS is treated as a tunable parameter. In our setup, entity-bearing nouns and relational terms are preserved, while adjectives and adverbs are randomly masked. This approach recognizes the importance of entities while introducing variation through modifiers.
Table C summarizes the performance on the RefCOCO validation set. All four strategies outperform the previous SoTA method, CARIS (CARIS) (74.67 oIoU), with consistent margins. Among them, WWM tends to remove entire entities, leading to slightly lower performance (76.03 oIoU). Span masking performs comparably to token-level masking (76.46 vs. 76.49). The POS-aware strategy achieves the best result (76.52), highlighting the importance of preserving semantic anchors while introducing controlled lexical noise. These findings, combined with our image masking study, demonstrate that carefully selecting what to mask in both visual and textual modalities is critical to achieving robust improvements.
Method val testA testB
DCL (Ours) 76.49 78.96 73.96
Cross-distillation 76.68 78.92 73.21
Table D: Impact of distillation strategy. The performance (oIoU) is evaluated across various distillation strategies on RefCOCO.
A.4 Analysis of Distillation Strategy
We further investigated whether incorporating a cross-modal contrastive loss (wang2024cm) could enhance our distortion-aware contextual learning (DCL) framework. As shown in Table D, adding cross-distillation yields a slight gain on the validation set (76.68 vs. 76.49), but this improvement does not generalize to the test splits. On RefCOCO testA, performance remains almost identical (78.92 vs. 78.96), and on testB it even decreases (73.21 vs. 73.96). In addition, this setting substantially increases computational cost, as each training step requires three forward passes instead of one. These results confirm that cross-distillation does not provide consistent accuracy benefits and results in an unfavorable trade-off between performance and efficiency.
A.5 Impact of Hyperparameters
ps 8 16 32 64 112
oIoU 76.58 76.49 76.49 76.47 76.01
Table E: Impact of patch size on image masking. The performance (oIoU) is evaluated across different patch sizes on the RefCOCO validation set. ps denotes the patch size.
$\lambda$ 0.1 0.25 0.5 0.75 1.0
oIoU 75.48 76.62 76.49 76.16 74.67
Table F: Impact of $\lambda$ on DCL procedure. The performance (oIoU) is evaluated across various $\lambda$ on the RefCOCO validation set.
Patch Size of Image Masking. To further analyze the hyperparameters of image masking, we examined the training results of different patch sizes for patch-wise image masking. Table E shows that our method consistently achieves superior performance regardless of patch size, outperforming the current SoTA method, CARIS (CARIS). Specifically, we observe that a patch size of 8 yields the highest performance. Nevertheless, the results suggest that our method exhibits stable and strong performance as long as the patch size remains within a reasonable range, highlighting its robustness to patch size variations.
Loss Weight of DCL. As described in Sec. 3.3, our training objective of Distortion-aware Contextual Learning (DCL) is $\mathcal{L}_{total}=\mathcal{L}_{dist}+\mathcal{L}_{ce}.$ To balance the scale of the loss values with previous methods (CARIS) during the training phase, we adapt the equation to:
$\mathcal{L}_{total}=(1-\lambda)\cdot\mathcal{L}_{dist}+\lambda\cdot\mathcal{L}_{ce},$ (5)
where $\lambda$ is set to 0.5 for all experiments to avoid extensive hyperparameter search for optimal performance. A $\lambda$ of 1.0 indicates training only with $\mathcal{L}_{ce}$, which is the equivalent to CARIS (CARIS). In Table F, we explore how varying $\lambda$ affects the DCL procedure. As shown in Table F, while the best performance is achieved with $\lambda=0.25$, the performance with $\lambda$ of 0.5 remains significantly better than the results without $\mathcal{L}_{dist}$. These results confirm that MaskRIS does not require extensive hyperparameter search for $\lambda$ and performs reasonably well with $\lambda$ of 0.5.